OpenStudy (anonymous):

Prove that 0.9999.... is the same as 1. =))

OpenStudy (asadkarim7):

if we round it off it would be same as 1

OpenStudy (anonymous):

lol, ok. That's not a proof though.

OpenStudy (asadkarim7):

lol why u want a proof , it can never be same as 1 until we round it off to avoid repetition

OpenStudy (anonymous):

I think she is referring to limit

OpenStudy (anonymous):

Yes it can. :/

OpenStudy (anonymous):

1/3 = .3333333..... .3333.... .3333.... +.3333... --------- .9999... So 1 = 1/3 + 1/3 + 1/3 = .9999...

OpenStudy (anonymous):

u hv received its ans in the other question.........

OpenStudy (anonymous):

Harkirat: I know. But I had already posted this before, wouldn't just ignore what people answered here.